OpenStudy (anonymous):

a family’s well pumps water into a tank at a rate of 10 gal/min. In the morning, the tank has 1700 gal. Which equation models the amount of water y in the tank after the well pumps for x min

OpenStudy (imstuck):

The equation would be y = 1700 + 10x, where y is the amount of water after all the pumping, 1700 is the initial start in the morning, and 10x is the rate of the water pumping (10 gallons) at x minutes. So if we look at x to be 15 minutes, after fifteen minutes with 1700 gallons to start we would have y = 1700 +10(15). This is basically f(15)=1700 + 10x. That's if we use 15 as an example. You could find the level of the water after however many minutes you want! Just fill in the number of minutes for x into the equation. Ok?

OpenStudy (anonymous):

thank you!

OpenStudy (imstuck):

You bet!